Suppose that Calloway golf would like to capitalize on Phil Michelson winning the Open Championship in 2013 by releasing a new putter. The new product will require new equipment for $420,831.00 that will be depreciated using the 5-year MACRS schedule. The project will run for 2 years with the following forecasted numbers:

Year 1 Year 2
Putter price $63.52 $63.52
Units sold 19,097.00 11,551.00
COGS 38.00% of sales 38.00% of sales
Selling and Administrative 18.00% of sales 18.00% of sales


Calloway has a 14.00% cost of capital and a 36.00% tax rate. The firm expects to sell the equipment after 2 years for a NSV of $141,417.00.

a) What is the project cash flow for year 1?

b) What is the project cash flow for year 2? (include the terminal cash flow here)

* use the equationn FCF1= (sales-expenses-depreciation(1-T) + depreciation - average net working capital - average gross PPE +- side effects
Answer #1
0 Year 1 Year 2
Putter price $               63.52 $         63.52
Units sold 19097 11551
Sales $       12,13,041 $   7,33,720
COGS (38% of sales) $         4,60,956 $   2,78,813
Selling and administrative (18% of sales) $         2,18,347 $   1,32,070
Depreciation (MACRS 5 Yr) $             84,166 $   1,34,666
NOI $         4,49,572 $   1,88,171
Tax at 36% $         1,61,846 $       67,741
NOPAT $         2,87,726 $   1,20,429
Add: Depreciation $             84,166 $   1,34,666
OCF $         3,71,892 $   2,55,095
Capital expenditure $      4,20,831
After tax salvage value [See workings below] $ -1,63,226
Project cash flow $    -4,20,831 $         3,71,892 $   4,18,322
Afer tax NSV:
NSV 141417
Book value = 420831*48% = 201999
Loss on sale 60582
Tax shield on loss at 36% 21809
Afer tax NSV = 141417+21809 = 163226